A steel beam is used in the construction of a skyscraper. By what fraction L/Lo does the length of the beam increase when the temperature changes from that on a cold winter day (–15 °F) to that on a summer day (+105 °F)?


1
Expert's answer
2021-08-29T16:56:04-0400

This ratio is


"\frac{L_S}{L_W}=\frac{1+\alpha T_S}{1+\alpha T_W},\\\space\\
\frac{L_S}{L_W}=\frac{1+0.0000061·105}{1+0.0000061 ·(-15)}=1.0007."